OpenStudy (anonymous):

help me solve this DE please. (2x+y)dx - (3x+2y-2)dy=0

OpenStudy (unklerhaukus):

\[(2x+y)\text dx - (3x+2y-2)\text dy=0\]\[(2x+y)\text dx = (3x+2y-2)\text dy\] \[\frac{2x+y}{3x+2y-2}~= \frac{\text dy}{\text dx}\]

OpenStudy (unklerhaukus):

im not sure there is a nice solution

hartnn (hartnn):

have u tried?till which step did u reach,@UnkleRhaukus gave u proper start... which substitution u have tried??

OpenStudy (anonymous):

u can show that the sub \(X=x−x_0\) and \(Y=y−y0\) will change the nonhomogeneous diff equation\[\frac{dy}{dx}=\frac{ax+by+c}{ex+fy+g}\]to homogeneous diff equation\[\frac{dY}{dX}=\frac{aX+bY}{eX+fY}\]where\( (x_0,y_0)\) is solution of system\[ax+by+c=0\]\[ex+fy+g=0\]

OpenStudy (anonymous):

non exact ode .i think integrating factor can be used to make it exact.

OpenStudy (anonymous):

and yes sami...also integrating factor will work

OpenStudy (anonymous):

ok. i'll try using integrating factor. thank you.

OpenStudy (anonymous):

let us know if its workin or not...

OpenStudy (anonymous):

do you think bernoulli's equation would help?

OpenStudy (anonymous):

emm...i dont think so

OpenStudy (anonymous):

it's hard, to get the integrating factor. lots of variables here.

OpenStudy (unklerhaukus):

\[\frac{\text dy}{\text dx}=\frac{2x+y}{3x+2y-2}\qquad\Longrightarrow\frac{\text dY}{\text dX}=\frac{2X+Y+2x_0+y_0}{3X+2Y+3x_0+2y_0-2}\] \[\qquad\qquad\qquad \qquad\qquad\qquad\qquad2x_0+y_0=0\]\[\qquad\qquad\qquad\qquad\qquad\qquad 3x_0+2y_0-2=0\]

OpenStudy (unklerhaukus):

\[3x_0+2(-2x_0)-2=0\]\[-x_0-2=0\]\[x_0=-2\] \[2(-2)+y_0=0\]\[y_0=4\]

OpenStudy (unklerhaukus):

\[\frac{\text dY}{\text dX}=\frac{2X+Y}{3X+2Y}\]\[\qquad=\frac{2+Y/X}{3+2Y/X}\] \[V=Y/X\qquad Y=VX\] \[\frac{\text d Y}{\text dX}=V+V'X\] \[V+V'X=\frac{2+V}{3+V}\] \[V'X=\frac{2+V}{3+V}~-V\] which is separable

OpenStudy (anonymous):

now how am i going to integrate this: (2z+1)dz / (z^2 - z -1)

OpenStudy (unklerhaukus):

\[\frac{2z+1}{z^2 - z -1}~=\frac{2z-1}{z^2 - z -1}~+\frac{2}{z^2-z-1}\]

OpenStudy (anonymous):

is this a partial fraction?

OpenStudy (unklerhaukus):

the first fraction has the numerator as the derivative for the denominator,

OpenStudy (anonymous):

yes of course! thanks a lot!

OpenStudy (anonymous):

how come the numerator of the first fraction becaME 2z-1?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

well we had to add 2 to get like that

OpenStudy (unklerhaukus):

\[\int\frac{2z+1}{z^2 - z -1}~~\text dz ~=\int\frac{2z-1}{z^2 - z -1}~~\text dz+\int\frac{2}{z^2-z-1}~~\text dz\]

OpenStudy (unklerhaukus):

the first intergal is easy now \[u= z^2-z-1\]\[\text du=(2z-1) \text dz\] \[=\int\frac{\text du}{u}+\int\frac{2}{z^2-z-1}~~\text dz\]

OpenStudy (anonymous):

so how about the next fraction? how am i going to integrate it?

OpenStudy (anonymous):

firstly complete the square for denum

OpenStudy (anonymous):

how about -5/4 after completing the square, where does it go?

OpenStudy (anonymous):

try to use this :\[\int \frac{u'}{u^2-a^2}dx = \frac{1}{2a}Log\left | \frac{u-a}{u+a} \right | + C\]

OpenStudy (anonymous):

sorry \[\int \frac{u'}{u^2-a^2}dx = \frac{1}{2a}\ln\left | \frac{u-a}{u+a} \right | + C\]

OpenStudy (anonymous):

ok. thank you. i have another DE to solve. help me again please.

OpenStudy (anonymous):

np post it...